Solve: (0≤x<2π)    a. tan 2x = cot 2x b. 2cos^2 x+cosx - 1=0

Solve: (0≤x<2π)

   a. tan 2x = cot 2x

b. 2cos^2 x+cosx - 1=0
Answer #1

(0) Tos given tanan lotan 102 XL2T) r we know that sinar wszu - losan sinar tano = sine coso colo = coso sino Sinaux sinau =외 ㅠ | + 1 For 야1C2T - 020 ) 5 = T42T M = A = = 3 N= + 3 , 3 8 수6T L IT CT ㅠ커니T 껸 | M = 5 - - T. 50 T T+ = ! - = 꺄6 So In integiven 26032x + L0SX-1 = 0 200sze + 26052-652l - l = 0 26 sul wsx +1) -I (OSX+1) = 0 . (2605x-1). I l0sx + 1) = 0 GOSH = , 6sx